# Long-range Meta-path Search on Large-scale Heterogeneous Graphs

### v4

@@ Abstract
Utilizing long-range dependency, though extensively studied in homogeneous graphs, has not been well investigated on heterogeneous graphs. Addressing this research gap presents two major challenges. The first is to alleviate computational costs while endeavoring to leverage as much effective information as possible in the presence of heterogeneity. The second involves overcoming the well-known over-smoothing issue occurring in various graph neural networks. To this end, we investigate the importance of different meta-paths and introduce an automatic framework for utilizing long-range dependency on heterogeneous graphs, denoted as Long-range Meta-path Search through Progressive Sampling (LMSPS). Specifically, we develop a search space with all meta-paths related to the target node type. By employing a progressive sampling algorithm, LMSPS dynamically shrinks the search space with hop-independent time complexity. Utilizing a sampling evaluation strategy as the guidance, LMSPS conducts a specialized and effective meta-path selection. Subsequently, only effective meta-paths are employed for retraining to reduce costs and overcome the over-smoothing issue. Extensive experiments on various heterogeneous datasets demonstrate that LMSPS discovers effective long-range meta-paths and outperforms the state-of-the-art. Besides, it ranks top-$1$ on the leaderboards of ogbn-mag in Open Graph Benchmark [^1] https://ogb.stanford.edu/docs/leader_nodeprop/#ogbn-mag. Our code is available at https://github.com/JHL-HUST/LDMLP.
Machine Learning, ICML
@@ 1 Introduction
Utilizing long-range dependency is essential for graph representation learning. For homogeneous graphs, many graph neural networks (GNNs) (Li et al., 2019; Bianchi et al., 2020; Alon & Yahav, 2021; Ying et al., 2021) have been developed to gain benefit from long-range dependency. Utilizing long-range dependency is also crucial for heterogeneous graphs. For instance, the Internet movie database (IMDB) contains $21K$ nodes with only $87K$ edges. Such sparsity means each node has only a few directly connected neighbors and requires models to enhance the node embedding from long-range neighbors. The main challenge in using long-range dependency on heterogeneous graphs is how to alleviate costs while striving to effectively utilize information in exponentially increased receptive fields, which is much more challenging compared to homogeneous graphs due to the heterogeneity. In addition, the well-known over-smoothing issue (Li et al., 2018; Keriven, 2022) occurring in many GNNs also needs to be addressed.
@@
Heterogeneous Graph Neural Networks (HGNNs) are popular deep learning techniques for heterogeneous graph representation learning. Their key idea is to aggregate valuable neighbor information based on a range of relations to enhance the semantics of vertices. Traditional HGNNs are typically classified into two categories: metapath-free methods and metapath-based methods. Most metapath-free HGNNs (Zhu et al., 2019; Hong et al., 2020; Hu et al., 2020b; Lv et al., 2021) utilize information from $l$-hop neighborhoods through stacking $l$ layers. However, on large-scale heterogeneous graphs, the number of nodes in receptive fields grows exponentially with the number of layers, making them hard to expand to large hops. A recent work (Mao et al., 2023) employs the graph transformer (Ying et al., 2021) to learn heterogeneous graphs, which can only exploit long-range dependency in small datasets due to the quadratic complexity of transformer (He et al., 2023).
@@
Metapath-based HGNNs (Wang et al., 2019; Fu et al., 2020; Ji et al., 2021; Yang et al., 2023) obtain information from $l$-hop neighborhoods by utilizing single-layer structures and meta-paths with maximum hop $l$, i.e., all meta-paths no more than $l$ hops. To exploit long-range dependency, the maximum hop needs to be large enough as there are no stacking layers. However, the number of meta-paths also grows exponentially with maximum hop $l$, corresponding to exponential receptive fields associated with layers in metapath-free methods. For instance, on ogbn-mag, to gain a $3$-hop meta-path based on the $2$-hop meta-path PAP, the next node type has three choices (A, P, and F). For each additional hop, the number of possible meta-paths increases exponentially. Hence, utilizing long-range dependencies on large-scale heterogeneous graphs has not been resolved yet.
@@
This paper investigates the importance of various meta-paths and makes two key observations: (1) A small number of meta-paths dominate the performance, and (2) certain meta-paths can have a negative impact on performance. The second observation explains why few HGNNs gain benefit from long-range neighbors. Different from homogeneous graphs, messages on heterogeneous graphs can be noisy or redundant for specific tasks, and the presence of long-range dependencies makes it more challenging to exclude negative heterogeneous information, even with the use of attention mechanisms. These findings highlight the opportunity to leverage long-range dependencies by selectively utilizing effective meta-paths.
@@
  • We propose a novel meta-path search framework termed LMSPS, that is the first HGNNs to utilize long-range dependency in large-scale heterogeneous graphs.
@@ 2 Preliminaries
Heterogeneous graph (Sun & Han, 2012). A heterogeneous graph is defined as $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{T},\mathcal{R},f_{\mathcal{T}},f_{\mathcal{R}}\}$ with $|\mathcal{T}{+}\mathcal{R}|{>}2$, where $\mathcal{V}$ denotes the set of nodes, $\mathcal{E}$ denotes the set of edges, $\mathcal{T}$ is the node-type set and $\mathcal{R}$ the edge-type set. Each node $v_{i}\in\mathcal{V}$ is maped to a node type $f_{\mathcal{T}}(v_{i})\in\mathcal{T}$ by mapping function $f_{\mathcal{T}}:\mathcal{V}\rightarrow\mathcal{T}$. Similarly, each edge $e_{t\leftarrow s}\in\mathcal{E}$ ($e_{ts}$ for short) is mapped to an edge type $f_{\mathcal{R}}(e_{ts})\in\mathcal{R}$ by mapping function $f_{\mathcal{R}}:\mathcal{E}\rightarrow\mathcal{R}$.
@@
Meta-path (Sun et al., 2011). A meta-path $P$ is a composite relation that consists of multiple edge types, i.e., $P\triangleq c_{1}\stackrel{{\scriptstyle r_{12}}}{{\longleftarrow}}c_{2}\cdots c_{l-1}\stackrel{{\scriptstyle r_{(l-1)l}}}{{\longleftarrow}}c_{l}$ ($P=c_{1}c_{2}\cdots c_{l}$ for short), where $c_{1},\dots,c_{l}\in\mathcal{T}$ and $r_{12},\dots,r_{(l-1)l}\in\mathcal{R}$.
@@
A meta-path $P$ corresponds to multiple meta-path instances in the underlying heterogeneous graph. For example, meta-path PAP corresponds to all paths of co-author relationships on the heterogeneous graph. In HGNNs, using meta-paths means selectively aggregating neighbors on the meta-path instances.
@@
3 Related Works
Heterogeneous Graph Neural Networks. HGNNs are proposed to learn rich and diverse semantic information on heterogeneous graphs. Several HGNNs (Li et al., 2021a; Ji et al., 2021; Lv et al., 2021; Mao et al., 2023) have involved high-order semantic aggregation. However, their methods are not applied to large-scale datasets due to relatively high costs. Additionally, many HGNNs (Yun et al., 2019; Hu et al., 2020b; Wang et al., 2019; Ji et al., 2021) have implicitly learned meta-paths by attention. However, few work employs the discovered meta-paths to generate final results, let alone generalize them to other HGNNs to demonstrate their effectiveness. For example, GTN (Yun et al., 2019) and HGT (Hu et al., 2020b) only list the discovered meta-paths. HAN (Wang et al., 2019), HPN (Ji et al., 2021) and MEGNN (Chang et al., 2022) validate the importance of discovered meta-paths by experiments not directly associated with the learning task. GraphMSE (Li et al., 2021b) is the only work that shows the performance of the discovered meta-paths. However, they are not as effective as the full meta-path set. Consequently, their learned meta-paths are not effective enough. Unlike these methods, the meta-paths searched by LMSPS not only enable it to outperform existing HGNNs but also are effective on other HGNNs. Furthermore, the meta-paths searched by LMSPS enable it to efficiently leverage long-range dependencies on large-scale datasets by only utilizing effective semantic information.
Meta-structure Search on Heterogeneous Graphs. Recently, some works have attempted to utilize neural architecture search (NAS) to discover meta-structures. GEMS (Han et al., 2020) is the first NAS method on heterogeneous graphs, which utilizes an evolutionary algorithm to search meta-graphs for recommendation tasks. DiffMG (Ding et al., 2021) searches for meta-graphs by a differentiable algorithm to conduct an efficient search. PMMM (Li et al., 2023) performs a stable search to find meaningful meta-multigraphs. However, meta-path-based HGNNs are mainstream methods (Schlichtkrull et al., 2018; Zhang et al., 2019; Wang et al., 2019; Fu et al., 2020), while meta-graph-based HGNNs are specialized. So, their searched meta-graphs are extremely difficult to generalize to other HGNNs. RL-HGNN (Zhong et al., 2020) proposes a reinforcement learning (RL)-based method to find meta-paths. On recommendation tasks, RMS-HRec (Ning et al., 2022) also proposes an RL-based meta-path selection strategy to discover meta-paths. However, how to search meta-structures on large-scale heterogeneous graphs is a non-trivial problem, and how to search meta-structures that are effective across different HGNNs is also challenging. All the above works have not addressed both problems.
Figure 2: The overall framework of LMSPS. Based on the progressive sampling and sampling evaluation in the search stage, the training stage employs $M$ effective meta-paths instead of the full $K$ target-node-related meta-paths. This figure exhibits aggregation of meta-paths with maximum hop $2$, i.e., 00, $1$, and $2$-hop meta-paths. The weight updates of feature projection are not shown for ease of illustration.
4 Motivation of Meta-path Search
Many HGNNs (Yun et al., 2019; Li et al., 2021b; Chang et al., 2022; Yang et al., 2023) employ all target-node-related meta-paths for heterogeneous graph representation learning. However, the contribution of each meta-path is under-explored. In this section, we conduct an in-depth study of different meta-paths and gain two crucial conclusions through experiments, helping us to propose the key idea of leveraging long-range dependencies effectively.
SeHGNN (Yang et al., 2023) employs attention mechanisms to fuse all the target-node-related meta-paths and outperforms the existing HGNNs. It has an important finding that models with single-layer structures and long meta-paths outperform those with multi-layers and short meta-paths, indicating the importance of long-range meta-paths. However, because the number of meta-paths increases exponentially with maximum hops, SeHGNN has to use a small maximum hop of $2$ for large-scale datasets like ogbn-mag (Hu et al., 2021) to save memory and reduce costs, which is insufficient as shown in Table 5 of the experiments.
[…]
In each experiment on DBLP, we remove one meta-path and compare the performance with the result of leveraging the full meta-path set to analyze the importance of the removed meta-path. As shown in Figure 1 (a), removing either A or AP or APA or APT has little impact on the final performance. However, removing APV results in severe degradation in performance, demonstrating APV is the critical meta-path on DBLP when $l=2$. We further retain one meta-path and remove others as shown in Figure 1 (b). The performance of utilizing APV is only slightly degraded compared to the full meta-paths set. Consequently, we obtain the first finding: A small number of meta-paths provide major contributions. In each experiment on ACM, we remove some meta-paths to analyze their impact on the final performance. Results in Figure 1 (c) show that the performance of SeHGNN improves after removing a part of meta-paths. For example, after removing PC and PCP, the Micro-F1 scores improve by $0.52\%$. So, we can conclude the second finding: Certain meta-paths can have a negative impact for heterogeneous graphs.
@@ 4 Motivation of Meta-path Search
The second finding is reasonable because heterogeneous information is not consistently beneficial for various tasks compared to homogeneous information. It is supported by the fact that various recent HGNNs (Ding et al., 2021; Li et al., 2023; Yang et al., 2023) have removed some edge types to exclude corresponding heterogeneous information during pre-processing based on substantial domain expertise or empirical findings. This observation explains why most HGNNs use a maximum hop of $2$, i.e., it is hard to exclude negative information under larger maximum hops. Additionally, because SeHGNN employs an attention mechanism, the performance degradation indicates the attention mechanism has limitations in dealing with noise.
@@
Motivated by the above findings, we can employ effective meta-paths instead of the full meta-path set without sacrificing performance. Although the number of meta-paths grows exponentially with the maximum hop, the proportion of effective meta-paths is small, which is similar to the $80/20$ rule of Pareto principle (Sanders, 1987; Erridge, 2006). To keep efficiency while trading off the performance, we choose a fixed number of meta-paths (like $30$) over all datasets, avoiding computation on numerous noisy or redundant meta-paths when the maximum hop is large.
@@ 5 The Proposed Method
The key point of our LMSPS is to utilize a search stage to reduce the exponentially increased meta-paths to a subset that is effective to the current dataset and task. It can overcome the main challenges of utilizing long-range dependency on heterogeneous graphs. First, utilizing effective meta-paths instead of all meta-paths alleviates computational and memory costs while keeping effective heterogeneous information. Second, each target node only aggregates neighbors on the path instances of effective meta-paths. Because the proportion of effective meta-paths is small and each node has a different neighborhood condition, each target node aggregates different neighbors under the constraints of effective meta-path instances. In this way, the over-smoothing issue can also be overcome.
@@
The super-net aims to automatically discover effective meta-paths for specific datasets or tasks, so the search results should not be affected by specific modules. Based on this consideration, we develop a simple MLP-based instead of transformer-based architecture for meta-path search because the former involves fewer human interventions (Tolstikhin et al., 2021; Wang et al., 2022). The super-net contains five blocks: neighbor aggregation, feature projection, progressive sampling search, sampling evaluation, and MLP.

### v6

@@ Abstract
Utilizing long-range dependency, a concept extensively studied in homogeneous graphs, remains underexplored in heterogeneous graphs, especially on large ones, posing two significant challenges: Reducing computational costs while maximizing effective information utilization in the presence of heterogeneity, and overcoming the over-smoothing issue in graph neural networks. To address this gap, we investigate the importance of different meta-paths and introduce an automatic framework for utilizing long-range dependency on heterogeneous graphs, denoted as Long-range Meta-path Search through Progressive Sampling (LMSPS). Specifically, we develop a search space with all meta-paths related to the target node type. By employing a progressive sampling algorithm, LMSPS dynamically shrinks the search space with hop-independent time complexity. Through a sampling evaluation strategy, LMSPS conducts a specialized and effective meta-path selection, leading to retraining with only effective meta-paths, thus mitigating costs and over-smoothing. Extensive experiments across diverse heterogeneous datasets validate LMSPS’s capability in discovering effective long-range meta-paths, surpassing state-of-the-art methods. Our code is available at https://anonymous.4open.science/r/LMSPS-BBEC.
@@ 1 Introduction
Utilizing long-range dependency is essential for graph representation learning. For homogeneous graphs, many graph neural networks (GNNs) [26, 3, 1, 53] have been developed to gain benefit from long-range dependency. Utilizing long-range dependency is also crucial for heterogeneous graphs. For instance, the Internet movie database (IMDB) contains $21K$ nodes with only $87K$ edges. Such sparsity means each node has only a few directly connected neighbors and requires models to enhance the node embedding from long-range neighbors. The main challenge in using long-range dependency on heterogeneous graphs is how to alleviate costs while striving to effectively utilize information in exponentially increased receptive fields, which is much more challenging compared to homogeneous graphs due to the heterogeneity. In addition, the well-known over-smoothing issue [28, 22] occurring in many GNNs also needs to be addressed.
@@
Heterogeneous Graph Neural Networks (HGNNs) are popular deep learning techniques for heterogeneous graph representation learning. The key idea is to aggregate valuable neighbor information based on a range of relations to enhance the semantics of vertices. Traditional HGNNs are typically classified into two categories: metapath-free methods and metapath-based methods. Most metapath-free HGNNs [62, 16, 19, 32] utilize information from $l$-hop neighborhoods by stacking $l$ layers. However, on large-scale heterogeneous graphs, the number of nodes in receptive fields grows exponentially with the number of layers, making them hard to expand to large hops. A recent work [34] employs the graph transformer [53] to learn heterogeneous graphs, which can only exploit long-range dependency in small datasets due to the quadratic complexity of transformer [15].
@@
Metapath-based HGNNs [46, 10, 21, 51] obtain information from $l$-hop neighborhoods by utilizing single-layer structures and meta-paths with maximum hop $l$, i.e., all meta-paths no more than $l$ hops. Benefiting from the feature of selective aggregation of neighbors based on meta-path, metapath-based HGNNs show greater potential in handling large-scale datasets [54, 51]. For metapath-based HGNNs, to exploit long-range dependency, the maximum hop needs to be large enough as there are no stacking layers. However, the number of meta-paths also grows exponentially with maximum hop $l$, corresponding to exponential receptive fields associated with layers in metapath-free methods. For instance, on OGBN-MAG, to gain a $3$-hop meta-path based on the $2$-hop meta-path PAP, the next node type has three choices (A, P, and F). For each additional hop, the number of possible meta-paths increases exponentially. Hence, utilizing long-range dependencies on large-scale heterogeneous graphs has not been resolved yet.
@@
This paper investigates the importance of various meta-paths and makes two key observations: (1) A small number of meta-paths dominate the performance, and (2) certain meta-paths can have a negative impact on performance. The second observation explains why few HGNNs gain benefit from long-range neighbors. Different from homogeneous graphs, messages on heterogeneous graphs can be noisy or redundant for specific tasks, and the presence of long-range dependencies makes it more challenging to exclude negative heterogeneous information, even with the use of attention mechanisms. For example, focusing too much on Institution instead of related Papers is harmful to predicting a paper’s field. These observations highlight the opportunity to leverage long-range dependencies by selectively utilizing effective meta-paths.
@@
  • We propose a novel meta-path search framework termed LMSPS, which to our knowledge is the first HGNNs to utilize long-range dependency in large-scale heterogeneous graphs.
@@ 2 Preliminaries
Heterogeneous graph [41]. A heterogeneous graph is defined as $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{T},\mathcal{R},f_{\mathcal{T}},f_{\mathcal{R}}\}$ with $|\mathcal{T}{+}\mathcal{R}|{>}2$, where $\mathcal{V}$ denotes the set of nodes, $\mathcal{E}$ denotes the set of edges, $\mathcal{T}$ is the node-type set and $\mathcal{R}$ the edge-type set. Each node $v_{i}\in\mathcal{V}$ is maped to a node type $f_{\mathcal{T}}(v_{i})\in\mathcal{T}$ by mapping function $f_{\mathcal{T}}:\mathcal{V}\rightarrow\mathcal{T}$. Similarly, each edge $e_{t\leftarrow s}\in\mathcal{E}$ ($e_{ts}$ for short) is mapped to an edge type $f_{\mathcal{R}}(e_{ts})\in\mathcal{R}$ by mapping function $f_{\mathcal{R}}:\mathcal{E}\rightarrow\mathcal{R}$.
@@
Meta-path [42]. A meta-path $P$ is a composite relation that consists of multiple edge types, i.e., $P\triangleq c_{1}\stackrel{{\scriptstyle r_{12}}}{{\longleftarrow}}c_{2}\cdots c_{l-1}\stackrel{{\scriptstyle r_{(l-1)l}}}{{\longleftarrow}}c_{l}$ ($P=c_{1}\cdots c_{l}$ for short), where $c_{1},\dots,c_{l}\in\mathcal{T}$ and $r_{12},\dots,r_{(l-1)l}\in\mathcal{R}$.
@@
A meta-path $P$ corresponds to multiple meta-path instances in the underlying heterogeneous graph. For example, meta-path APA corresponds to all paths of co-author relationships on the heterogeneous graph. Using meta-paths means selectively aggregating neighbors on the meta-path instances.
3 Related Works
Heterogeneous Graph Neural Networks. HGNNs are proposed to learn rich and diverse semantic information on heterogeneous graphs. Several HGNNs [27, 21, 32, 34] have involved high-order semantic aggregation. However, their methods are not applied to large-scale datasets due to high costs. Additionally, many HGNNs [55, 19, 46, 21] have implicitly learned meta-paths by attention. However, few work employs the discovered meta-paths to produce final results, let alone generalize them to other HGNNs to demonstrate their effectiveness. For example, GTN [55] and HGT [19] only list the discovered meta-paths. HAN [46], HPN [21] and MEGNN [4] validate the importance of discovered meta-paths by experiments not directly associated with the learning task. GraphMSE [29] is the only work that shows the performance of the discovered meta-paths. However, they are not as effective as the full meta-path set. So, their learned meta-paths are not effective enough.
Meta-structure Search on Heterogeneous Graphs. Recently, some works have attempted to utilize neural architecture search (NAS) to discover meta-structures. GEMS [14] is the first NAS method on heterogeneous graphs, which utilizes an evolutionary algorithm to search meta-graphs for recommendation tasks. DiffMG [7] searches for meta-graphs by a differentiable algorithm to conduct an efficient search. PMMM [25] performs a stable search to find meaningful meta-multigraphs. However, meta-path-based HGNNs are mainstream methods [39, 56, 46, 10], while meta-graph-based HGNNs are specialized. So, their searched meta-graphs are extremely difficult to generalize to other HGNNs. RL-HGNN [59] proposes a reinforcement learning (RL)-based method to find meta-paths. On recommendation tasks, RMS-HRec [35] also proposes an RL-based meta-path selection strategy to discover meta-paths. Both of them are very time-consuming.
4 Motivation of Long-range Meta-path Search
SeHGNN [51] employs attention mechanisms to fuse all the target-node-related meta-paths and outperforms the existing HGNNs. It has an important observation that models with single-layer structures and long meta-paths outperform those with multi-layers and short meta-paths, indicating the advantages of long-range meta-paths. However, because the number of meta-paths increases exponentially with maximum hops, SeHGNN has to use a small maximum hop to save memory and reduce costs. For example, The maximum hop is $2$ for large-scale datasets OGBN-MAG [18], which is insufficient, as shown in Table 3 of the experiments. This inspires us to consider whether we can only use effective meta-paths instead of all meta-paths to reduce the consumption of large maximum hops.
[…]
In each experiment on DBLP, we remove one meta-path and compare the performance with the result of leveraging the full meta-path set to analyze the importance of the removed meta-path. As shown in Figure 1 (a), removing either A or AP or APA or APT has little impact on the final performance. However, removing APV results in severe degradation in performance, demonstrating APV is the critical meta-path on DBLP when $l=2$. We further retain one meta-path and remove others as shown in Figure 1 (b). The performance of utilizing APV is only slightly degraded compared to the full meta-paths set. Consequently, we obtain the first observation: A small number of meta-paths provide major contributions. In each experiment on ACM, we remove some meta-paths to analyze their impact on the final performance. Results in Figure 1 (c) show that the performance of SeHGNN improves after removing a part of meta-paths. For example, after removing PC and PCP, the Micro-F1 scores are improved by $0.52\%$. So, we can conclude the second observation: Certain meta-paths can have a negative impact for heterogeneous graphs.
@@ 4 Motivation of Long-range Meta-path Search
The second observation is reasonable because heterogeneous information is not consistently beneficial for various tasks compared to homogeneous information. It is supported by the fact that various recent HGNNs [7, 25, 51] have removed some edge types to exclude corresponding heterogeneous information during pre-processing based on substantial domain expertise or empirical observations. This observation explains why most HGNNs use a maximum hop of $2$, i.e., it is hard to exclude negative information under larger maximum hops. Additionally, because SeHGNN employs an attention mechanism, the performance degradation indicates the attention mechanism has limitations in dealing with noise.
@@
Although long-range meta-paths outperform short meta-paths [51], they need a large maximum hop, resulting in exponentially increasing meta-paths. Motivated by the above two observations, unlike existing methods [4, 29, 51, 55], we can employ effective meta-paths instead of the full meta-path set without sacrificing performance. Although the number of meta-paths grows exponentially with the maximum hop, the proportion of effective meta-paths is small, which is similar to the $80/20$ rule of Pareto principle [38, 9]. To keep efficiency while trading off the performance, we choose a fixed number of meta-paths (like $30$) over all datasets. Therefore, the exponential meta-paths can be reduced to a constant. Now the key point is how to find effective meta-paths.
@@ 5 The Proposed Method
The key of our LMSPS is to utilize a search stage to reduce the exponentially increased meta-paths to a subset effective to the current dataset and task. It can overcome the main challenges of utilizing long-range dependency on heterogeneous graphs. First, utilizing effective meta-paths instead of all meta-paths alleviates computational and memory costs while keeping effective heterogeneous information. Second, each target node only aggregates neighbors on the path instances of effective meta-paths. Because the proportion of effective meta-paths is small and each node has a different neighborhood condition, each target node aggregates different neighbors under the constraints of effective meta-path instances. In this way, the over-smoothing issue can also be overcome.
@@
The super-net aims to automatically discover effective meta-paths for specific datasets or tasks, so the search results should not be affected by specific modules. Based on this consideration, we develop a simple MLP-based instead of transformer-based architecture for meta-path search because the former involves fewer human interventions [43, 47]. The super-net contains five blocks: neighbor aggregation, feature projection, progressive sampling search, sampling evaluation, and MLP.
Figure 2: The overall framework of LMSPS. Based on the progressive sampling and sampling evaluation in the search stage, the training stage employs $M$ effective meta-paths instead of the full $K$ target-node-related meta-paths. It exhibits aggregation of meta-paths with maximum hop $2$, i.e., 00, $1$, and $2$-hop meta-paths. The weight updates of feature projection are not shown for ease of illustration.